OpenStudy (anonymous):

A farmer builds a rectangular pen with a perimeter of 116 feet. If the length is 5 feet more than the width, what is the width?

OpenStudy (midhun.madhu1987):

Let the width be x and length = x + 5 Perimeter. P = 2(l + w) 2(x+ 5 + x) = 116

OpenStudy (midhun.madhu1987):

Solve for x.

OpenStudy (anonymous):

The width is 26.5 ft.